# A Simple DropConnect Approach to Transfer-based Targeted Attack

### v1

@@ Abstract
We study the problem of transfer-based black-box attack, where adversarial samples generated using a single surrogate model are directly applied to target models. Compared with untargeted attacks, existing methods still have lower Attack Success Rates (ASRs) in the targeted setting, i.e., the obtained adversarial examples often overfit the surrogate model but fail to mislead other models. In this paper, we hypothesize that the pixels or features in these adversarial examples collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation. Then, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD) to enhance transferability, by creating diverse variants of surrogate model at each optimization iteration. We conduct extensive experiments across various CNN- and Transformer-based models to demonstrate the effectiveness of MCD. In the challenging scenario of transferring from a CNN-based model to Transformer-based models, MCD achieves 13% higher average ASRs compared with state-of-the-art baselines. MCD  boosts the performance of self-ensemble methods by bringing in more diversification across the variants while reserving sufficient semantic information for each variant. In addition, MCD attains the highest performance gain when scaling the compute of crafting adversarial examples.
footnotetext: Preprint. Work in progress.
@@ 1 Introduction
In recent years, deep neural networks have become prevalent in computer vision applications [15, 17, 9], but are highly vulnerable to maliciously crafted inputs, which are called adversarial examples [34, 11]. A major concern is their transferability, that is, adversarial examples generated using a white-box model can be directly used to fool other black-box models, without any access to their architectures, parameters or gradients [13]. Since this type of attacks, usually referred to as transfer-based attacks, do not require any interaction with the target model, they pose severe security risks to real-world machine learning systems. Therefore, studying effective transfer-based attack methods is crucial to understanding the vulnerabilities, and in turn enhancing the robustness of machine learning models.
@@
This paper focuses on targeted transfer-based attacks with a single surrogate model, where the goal is to deceive black-box models to classify input images into a specific incorrect category. Due to the high complexity of decision boundaries, existing methods still have noticeably lower Attack Success Rates (ASRs) in the targeted setting than in the untargeted [47, 2]. A key observation is that the generated adversarial examples tend to overfit the surrogate model but fail to generalize to other models. To improve transferability, various strategies have since been proposed. Noticing that multiple surrogate models can help enhance transferability but in practice finding proper models for the same task is not easy [7, 26, 25, 46], model self-ensemble [18, 45, 22, 28] tries to create multiple models from only one accessible model, which is in accordance with our setting. Input transformation [44, 48, 1, 42, 40, 2] applies different transformations to inputs and diversify input patterns in order to reduce overfitting. A notable method is Clean Feature Mixup (CFM) [2], which randomly mixes high-level features with shuffled clean features and has achieved state-of-the-art performance. Despite these progresses, there is still much room for further improvement.
In this work, we hypothesize that there is a key aspect underlying the present overfitting issue: the pixels in many adversarial examples, obtained from existing transferred-based methods, collaborate in a highly dependent manner to maximize the success of adversarial attacks on the surrogate model, which we refer to as perturbation co-adaptation. Similar to the co-adaption issue of units in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that can fool the surrogate model but fail on other target models. How to further alleviate this co-adaptation issue would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
To this end, we propose to Mitigate perturbation Co-adaptation by DropConnect (MCD), motivated by the classic Dropout technique that can effectively prevent units from co-adaptation during training neural networks [39]. Rather than directly applying Dropout to the input or the perturbation, our method extends random masking to 1) high-level features that contain richer semantic information, and 2) model parameter space for more flexible and diverse feature disturbances for perturbation. Fig. 1 illustrates the proposed method. At each optimization iteration, MCD  creates diverse variants of the surrogate model with masked parameters, which are then used to update the adversarial examples. In this sense, the final algorithm can be regarded as a self-ensemble method. Compared with other related self-ensemble methods [45, 22], MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, hence improves the transfer attack performance. In addition, MCD can be easily implemented into all layers in the surrogate model, with little comprise in terms of attack performance. A more concrete comparison is left to Sec. 4.3.
@@
Figure 1: An illustration of the proposed method MCD. We apply MCD to selected layers in the surrogate model to create multiple and diverse variants at each optimization iteration.
@@
We evaluate the proposed method across various CNN- and Transformer-based target models, as well as against several defense methods. The experimental results show that MCD outperforms other state-of-the-art methods. In particular, in the challenging scenario of transferring from CNN-based model to Transformer-based models, MCD achieves 13.0% and 17.9% higher average ASRs with ResNet-50 [15] and DenseNet-121 [17] as surrogate models, respectively. Moreover, MCD achieves the highest performance gain when scaling the compute for crafting adversarial examples. Specifically, with ResNet-50 as surrogate model, increasing the optimization iterations from 300 to 500 and number of forward-backward passes per iteration from 1 to 5 boosts the average ASR by 15.9%.
@@
  • We show the perturbation co-adaptation issue in existing transfer-based attack methods in the targeted setting. Reducing this co-adaptation issue would in turn enhance transferability of adversarial examples.
  • We propose MCD to improve the transferability, by creating diverse variants of surrogate model during optimization. MCD brings in more diversification across the variants while reserving sufficient semantic information for each variant, and is also simple to implement.
  • Extensive experiments across various surrogate and target models validate the superior performance of MCD over other methods. Moreover, MCD benefits the most when we scale the compute of crafting adversarial examples.
2 Related Works
In this section, we provide a brief review of related works focusing on transfer-based attack methods.
The first class applies transformations to the input images, thereby diversifying the input patterns [44, 48, 8, 1, 42, 40, 2]. Diverse Inputs (DI) [44] is one of the representative methods. At each iteration, it applies random and differentiable transformations (e.g., random resizing) to the input image and then maximizes the loss function w.r.t. the transformed inputs. Resized Diverse Inputs (RDI) [48] extends DI by resizing the transformed image back to its original size. The Translation-Invariant (TI) attack method [8] optimizes perturbations across a set of translated images, making the adversarial example less sensitive to the surrogate model. Block Shuffle and Rotate (BSR) [40] divides the input image into blocks, followed by shuffling or rotation, and Object-based Diversity Inputs (ODI) [1] generates an adversarial example on a 3D object and leads the rendered image to being classified as the target class. Admix [42] mixes the input image with random samples from other classes. Clean Feature Mixup (CFM) [2] extends Admix to high-level feature space and introduces two types of competing noises to guide adversarial perturbations, achieving state-of-the-art performance in transfer-based targeted attacks.
Closely related to the present work is self-ensemble [18, 45, 22, 28], which creates multiple models from only one surrogate model. Transfer-based self-ensemble (T-sea) [18] targets the task of object detection, and the self-ensemble method in [28] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper. Ghost Network [22] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration. Masking Unimportant Parameters (MUP) [45] drops out unimportant parameters according to a predefined Taylor expansion-based metric.
@@ 3 Preliminary
Under the white-box setting where the model is fully accessible, adversarial example can be obtained through solving the following problem:
@@
where $\mathcal{L}(\cdot)$ represents a loss function like cross-entropy loss. A strong and popular attack method to the above optimization problem is the Iterative Fast Gradient Sign Method (I-FGSM) [20] that leverages the sign of the gradient information. In particular, at each iteration $t$, the adversarial example is updated by adding a small perturbation:
@@
In the black-box settings, the gradient information is not available and we only have limited access to the target model. A large class of attack methods in this setting is transfer-based, which assumes that the adversarial examples exhibit transferability across different neural models; that is, adversarial examples generated on one model to mislead not only that model but also other models [34]. Thus, one can use the above updates (2) on an available white-box model, usually called surrogate model, to generate adversarial examples that are applied to the target model. To boost the performance of transfer-based methods, an effective approach is to perform some transformation $\mathcal{T}$ (like resizing [44, 48] and scaling [24]) to the input and then use $\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$ as the gradient information.
@@ 4.1 Motivation
We hypothesize that the pixels or features in the adversarial example, obtained through solving Problem (1), collaborate in a highly dependent manner to maximize the success of an adversarial attack on the surrogate model, which we refer to as perturbation co-adaptation. Even with state-of-the-art transfer-based attack methods, this issue is still a key aspect that can make the perturbation fail to mislead other models.
We now conduct a pilot study to verify our hypothesis. Since adversarial example is a composition of the given clean image and some perturbation, Problem (1) is equivalent to finding a perturbation that maximizes the loss function. In this sense, we perform random masking with different Dropout ratios to the obtained perturbation $\delta:=x_{\mathrm{adv}}-x$, to disrupt potential co-adaptations among pixels. Note that the masked version of perturbation always satisfies the $\ell_{p}$-norm constraint if the adversarial example already meets this constraint. We consider a targeted attack setting on an ImageNet-compatible dataset (see more details in Sec. 5.1), where adversarial examples are crafted using ResNet-50. The used attack methods include DI, ODI and CFM [44, 1, 2].
The results are shown in Fig. 2 (additional results can be found in Sec. B.1). We can observe an obvious decrease in ASRs as the masking ratio exceeds 20%, in the DI, ODI and CFM methods, indicating that the pixels in the obtained perturbation are likely to be heavily dependent to attain a successful targeted attack. Similar to the co-adaption issue in training neural networks [33], we believe that perturbation co-adaptation would lead to overly complex patterns that mislead surrogate model but fail on other models. As such, how to further alleviate this co-adaptation would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
4.2 Mitigating Perturbation Co-adaptation by DropConnect
As per the pilot study, a direct approach to improving transferability would be to apply random masking to pixels in perturbation during the iterative updates. As shown in Sec. B.2, by encouraging perturbations to explore more diverse feature disturbances, this way indeed yields higher ASRs than the baseline DI and RDI methods in the targeted attack setting. However, it does not consistently achieve the best attack performance; a potential reason is that pixel-level masking may be less effective of capturing semantic information for image classification. Based on these initial results, it is tempting to extend this direct idea to craft more transferable adversarial examples.
[…]
(b) ResNet-50 $\to$ MobileNet-v2
Figure 2: ASRs(%) with increasing mask ratio.
The rest of this work is devoted to further mitigating the perturbation co-adaptation, by extending random masking to higher-level representations. The overall approach is to create multiple diverse variants of the surrogate model during optimization. We first consider the following questions:
  • Q1: how to achieve more flexible and diverse feature disturbances for perturbation than vanilla Dropout?
  • Q2: which intermediate representations or layers to perform the random masking?
For Q1, we observe that in typical machine learning problems where the goal is to optimize model parameters w.r.t. given training data, standard Dropout [33] is applied to the units or certain features. Finding adversarial example, as formulated in Problem (1), can be treated as optimizing perturbed image $x_{\mathrm{adv}}$ with fixed model parameters as ‘training data’ or ‘input features’. As analogy to Dropout, we can randomly drop out the model parameters connecting the units. This is indeed the DropConnect technique [39], a generalized version of Dropout by providing a more fine-grained way of regularizing the network during training. As such, we will focus on DropConnect in this paper.
@@ 4.2 Mitigating Perturbation Co-adaptation by DropConnect
For Q2, we notice that DropConnect is mainly effective in terms of linear layers [39]. We thus apply DropConnect to the weight and bias parameters of linear layers as well as the transformation parameters of normalization layers in the surrogate model. Both types of layers are widely used in mainstream architectures including Transformer [38]. Empirically, our ablation study in Sec. 5 validates the effectiveness of this choice, compared with convolutional layers.
@@
We now present our attack method, Mitigating perturbation Co-adaptation by DropConnect (MCD), as summarized in Algorithm 1. Take linear layer for example. We perform independent random masking onto the weight and bias (when present) parameters using Bernoulli sampling. Specifically, for surrogate model $f$, let $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denote the weight matrix and $b\in\mathbb{R}^{d_{\text{out}}}$ the bias vector associated with a selected linear layer. Here $d_{\text{in}}$ and $d_{\text{out}}$ represent the input and output dimensions, respectively. The corresponding masks for the weight matrix and bias are
$\mathcal{M}_{w}\sim\text{Bernoulli}(p_{w}),\quad\mathcal{M}_{b}\sim\text{Bernoulli}(p_{b}),$ (3)
@@
$W_{\mathcal{M}}=\mathcal{M}_{w}\odot W,\quad b_{\mathcal{M}}=\mathcal{M}_{b}\odot b,$ (4)
@@
Algorithm 1 Mitigating perturbation Co-adaptation by DropConnect (MCD)
@@
6:         Obtain modified model $f_{\mathcal{M}}$ by applying MCD to each layer in $\mathbb{L}$ according to Eqs. (3) and (4).
@@
Unlike standard Dropout or DropConnect in training neural networks, MCD does not need additional re-scaling. The random masks $\mathcal{M}_{w}$ and $\mathcal{M}_{b}$ in Eq. 3 are re-generated for each inference, producing diverse variants of the surrogate model with different parameters if we conduct multiple inferences at an iteration. In addition, MCD  can be naturally integrated with existing input transformation and gradient stabilization methods for crafting adversarial examples, as shown in Lines 5 and 8 in Algorithm 1.

### v2

# RaPA: Enhancing Transferable Targeted Attacks via Random Parameter Pruning

@@ Abstract
Compared to untargeted attacks, targeted transfer-based attack is still suffering from much lower Attack Success Rates (ASRs), although significant improvements have been achieved by kinds of methods, such as diversifying input, stabilizing the gradient, and re-training surrogate models. In this paper, we find that adversarial examples generated by existing methods rely heavily on a small subset of surrogate model parameters, which in turn limits their transferability to unseen target models. Inspired by this, we propose the Random Parameter Pruning Attack (RaPA), which introduces parameter-level randomization during the attack process. At each optimization step, RaPA randomly prunes model parameters to generate diverse yet semantically consistent surrogate variants.We show this parameter-level randomization is equivalent to adding an importance-equalization regularizer, thereby alleviating the over-reliance issue. Extensive experiments across both CNN and Transformer architectures demonstrate that RaPA substantially enhances transferability. In the challenging case of transferring from CNN-based to Transformer-based models, RaPA achieves up to 11.7% higher average ASRs than state-of-the-art baselines(with 33.3% ASRs), while being training-free, cross-architecture efficient, and easily integrated into existing attack frameworks.
@@ 1 Introduction
Deep neural networks have become prevalent in computer vision applications [16, 18, 10], but are highly vulnerable to maliciously crafted inputs, called adversarial examples [41, 12]. A major concern is their transferability, that is, adversarial examples generated using a white-box model can directly fool other black-box models, without any access to their architectures, parameters or gradients [14]. Since this type of attacks, usually referred to as transfer-based attacks, do not require any interaction with the target model, they pose severe security risks to real-world machine learning systems. Therefore, studying effective transfer-based attack methods is crucial to understand the vulnerabilities and further enhance model robustness.
@@
This paper focuses on targeted transfer-based attacks with a single surrogate model, where the goal is to deceive black-box models to classify input images into a specific incorrect category. Due to the high complexity of decision boundaries, existing methods still have noticeably lower Attack Success Rates (ASRs) in the targeted setting than in the untargeted [56, 2]. A key observation is that the generated adversarial examples tend to overfit the surrogate model but fail to generalize to other models. To improve transferability, various strategies have been proposed. Observing that multiple surrogate models can help enhance transferability but in practice finding proper models for the same task is not easy [8, 30, 29, 55], model self-ensemble [19, 54, 25, 33] tries to create multiple models from an accessible model. Input transformation [52, 57, 1, 48, 46, 49, 2, 26] applies different transformations to inputs and diversify input patterns to reduce overfitting. A notable method is Clean Feature Mixup (CFM) [2], which randomly mixes high-level features with shuffled clean features. Building upon it, Feature Tuning Mixup (FTM) [26] introduces learnable and attack-specific feature perturbations, achieving new state-of-the-art performance in transferability. Despite these progresses, there is still much room for further improvement.
In this work, we take a different perspective and identify a previously overlooked cause of poor transferability: the generated adversarial perturbations rely excessively on a small subset of parameters in the surrogate model, which limits their generalization to other models that have different parameter configurations or training dynamics. In other words, adversarial perturbation in existing methods tends to exploit a few “shortcut” parameters, leading to strong white-box performance but poor black-box transferability.
To mitigate this issue, we propose Random Parameter Pruning Attack (RaPA) that introduces parameter-level randomization into the attack process. At each optimization step, RaPA randomly prunes a subset of parameters in the surrogate model and uses multiple masked variants to update the adversarial example. We show that taking the expectation over such random masks is equivalent to adding an importance regularization term that aims to equalize parameter contributions, thus preventing over-reliance on a few dominant parameters. Conceptually, RaPA can be viewed as a self-ensemble method: each randomly pruned model represents a diverse yet semantically consistent variant of the surrogate. Previous self-ensemble approaches SASD-WS [50], MUP [54], and Ghost Network [25] rely on training-based model enhancement, deterministic pruning metrics, and structural perturbations, respectively. In contrast, RaPA is training-free, cross-architecture efficient, and straightforward to implement.
@@
Figure 1: An illustration of the proposed method RaPA. We apply RaPA to selected layers in the surrogate model to create multiple and diverse variants at each iteration.
@@
We evaluate the proposed method across various CNN- and Transformer-based target models, as well as against several defense methods. The experimental results show that RaPA outperforms other state-of-the-art methods. In particular, in the challenging scenario of transferring from CNN-based model to Transformer-based models, RaPA achieves 11.7% and 17.5% higher average ASRs with ResNet-50 [16] and DenseNet-121 [18] as surrogate models, respectively. Moreover, RaPA achieves the highest performance gain when scaling the compute for crafting adversarial examples. Specifically, with ResNet-50 as surrogate model, increasing the optimization iterations from 300 to 500 and number of forward-backward passes per iteration from 1 to 5 boosts the average ASR by 15.9%.
@@
  • We show that adversarial examples from existing transfer-based attacks rely heavily on a tiny subset of parameters in the surrogate model. Alleviating this over-reliance can in turn enhance the transferability of attack.
  • We propose the RaPA, which introduces parameter-level randomization during attack optimization. We show, both intuitively and empirically, that random pruning implicitly equalizes parameter importance, acting as a regularizer to mitigate the over-reliance issue.
  • Experiments across diverse surrogate and target models demonstrate that RaPA consistently outperforms existing methods. RaPA further benefits from increased computational budget, achieving larger improvements when scaling optimization iterations or inference steps.
Method DI RDI SI Admix ODI BSR CFM RaPA
No pruning 98.2 98.7 98.7 98.1 98.9 98.5 98.0 98.2
Pruning bottom 0.5% 98.2 98.7 98.7 98.0 98.9 98.5 98.0 98.0
Pruning top 0.5% 16.0 28.6 31.9 29.9 19.9 36.7 51.3 64.5
Table 1: ASRs (%) before and after pruning the selected subsets of parameters in the surrogate model on the ImageNet-compatible dataset. Detailed experimental setting can be found in Section 4.1.
@@ 2 Preliminary
This section introduces the background of adversarial attacks and briefly reviews related works on targeted transfer-based attacks. See Appendix A for more related work.
2.1 Background
@@
In the white-box setting where the model is fully accessible, adversarial example can be obtained by the following:
@@
where $\mathcal{L}(\cdot)$ represents a loss function like cross-entropy loss. The Fast Gradient Sign Method (FGSM) [12] uses the gradient direction to solve this problem and craft adversarial examples, while Iterative FGSM (I-FGSM) [22] extends this idea to an iterative scheme. In particular, at each iteration $t$, the adversarial example is updated by adding a small perturbation:
@@
2.2 Related Work
In black-box settings, the gradient information is not available and we only have limited access to the target model. Transfer-based method assumes that the adversarial examples generated on one model to mislead not only that model but also other models [41].
To improve transferability, many methods have been proposed. The first class, input-transformation techniques, applies a transformation $\mathcal{T}$ and uses $\nabla_{x_{\mathrm{adv}}^{t-1}}\mathcal{L}\left(f(\mathcal{T}(x_{\mathrm{adv}}^{t-1}))\right)$ as the gradient. Diverse Inputs (DI) [52] and its variant Resized DI (RDI) [57] apply random transformations to increase input variation during optimization. Translation-Invariant (TI) [9] averages gradients over translated inputs to reduce location sensitivity. Structure Invariant Attack (SIA) [49] and Block Shuffle and Rotate (BSR) [46] both perform block-level local transformations, with SIA applying diverse transforms and BSR focusing shuffle and rotate operations. Object-based DI (ODI) [1] generates adversarial examples rendered on 3D objects, while Admix [48] mixes inputs with random samples from other classes. CFM [2] extends it to the feature space with competing noises, while FTM [26] further adds learnable, attack-specific perturbations to achieve state-of-the-art transferability.
The second class focuses on stabilizing gradient updates to improve transferability. Momentum Iterative FGSM (MI-FGSM) [8] incorporates a momentum term into I-FGSM to help avoid local optima. Scale-Invariant (SI) optimization [28] improves transferability by applying perturbations across multiple scaled copies of the input, leveraging the scale-invariance property of deep models.
Beyond the above methods, another type of approaches re-train the surrogate model to enhance transferability, e.g., DSM [53] and SASD-WS [50] improve model generalization and transferability through knowledge distillation or sharpness-aware self-distillation.
Closely related to the present work is self-ensemble [19, 54, 25, 33], which creates multiple models from only one surrogate model. The self-ensemble method in [33] specifically considers vision Transformer as surrogate model and is denoted as SE-ViT in this paper. Ghost Network [25] perturbs surrogate model to create a set of new models and then samples one model from the set at each iteration. Masking Unimportant Parameters (MUP) [54] drops out unimportant parameters according to a predefined Taylor expansion-based metric, while Diversity Weight Purning(DWP) [45] only prunes the parameters with small absolute values.
@@ 3.1 Motivation
We observe that the adversarial perturbations generated by solving Problem (1) tend to rely heavily on a small subset of parameters in the surrogate model. These parameters may stem from specific training schemes, datasets, or architectural choices. As a result, adversarial examples that strongly depend on these parameters often fail to generalize and mislead other models. Even with SoTA transfer-based attack methods, this issue remains a key factor that can lead to the failure of adversarial example transfer.
To quantify the phenomenon, we conduct a pilot study using the framework of Optimal Brain Damage (OBD) [23, 32], which quantifies the importance of each model parameter from the perspective of sensitivity analysis. Specifically, given an adversarial example $x_{\mathrm{adv}}$ and a loss function $\mathcal{L}$. Let $\theta$} represent the entire set of model parameters. The importance of a parameter $\theta_{i}$ in the surrogate model $f$ is computed as:
$\mathcal{I}(\theta_{i})=\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}})}{\partial\theta_{i}^{2}}\times\theta_{i}^{2}.$ (3)
This metric reflects how much the loss would change if a parameter $\theta_{i}$ were removed, and can serve as a proxy for its contribution to the effectiveness of the adversarial example.
Next, we consider pruning two distinct subsets of the surrogate model’s parameters based on this importance metric: the top 0.5% most important and the bottom 0.5% least important parameters(see Section B.1 for details). For each adversarial example, we instantiate the model, prune the selected subset, and evaluate the ASR on the resulting model. Table 1 reports the ASRs after pruning the two subsets of model parameters. Here we use ResNet50 as the surrogate model and detailed setting can be found in Section 4.1.
We observe that pruning the most important parameters leads to a drastic drop in ASR—more than 46%, whereas pruning the least important parameters yields negligible impact. This observation suggests that adversarial examples generated by existing methods are highly dependent on the most important parameters, validating our observation on the over-reliance issue. As such, how to further alleviate this strong dependence on specific parameters would be a key to improving the transferability of adversarial examples over existing transfer-based attack methods.
3.2 Alleviating Over-reliance via Random Parameter Pruning
As per the pilot study, a direct approach to improving transferability would be to mask the most important parameters at each optimization step, thereby mitigating the over-dependency on them. However, accurately identifying important parameters requires computing second-order derivatives, which is computationally expensive for all parameters. Although we can approximate them with first-order terms, masking the most important parameters typically causes the surrogate model’s capacity to degrade rapidly, and the resulting adversarial examples may fail to fool the target model—and even the original surrogate model itself (See Section D.1 for further explanations). To address this problem, we propose to apply random parameter pruning to the surrogate model at each optimize step. This approach avoids expensive computations while achieving the goal of reducing over-reliance on specific parameters.
Intuition and Theoretical Explanation
Our core idea is that randomly pruning parameters at different optimization steps encourages the generated adversarial examples to be less dependent on particular parameter subsets. This in turn improves transferability across different target models.
We define a random binary mask $\mathcal{M}\in\{0,1\}^{|\theta|}$, where each entry is independently sampled from a Bernoulli distribution: $\mathcal{M}_{i}\sim\text{Bernoulli}(1-p)$. Here $p\in[0,1]$ is the probability of masking a parameter. With a small $p$, we would have $\mathbb{E}[\mathcal{M}_{i}]\approx 1$. Then the parameter of the model used in the forward pass becomes $\mathcal{M}\odot\theta$, where $\odot$ denotes element-wise multiplication.
Under this setup, the expected loss over random masks can be approximated using a second-order Taylor expansion:
$~\mathbb{E}_{\mathcal{M}}[\mathcal{L}(f(x_{\mathrm{adv}};\mathcal{M}\odot\theta))]$
$\approx$ $~\mathcal{L}(f(x_{\mathrm{adv}};\theta))+\frac{p(1-p)}{2}\sum_{i}\frac{\partial^{2}\mathcal{L}f(x_{\mathrm{adv}};\theta)}{\partial\theta_{i}^{2}}\theta_{i}^{2},$ (4)
which is sum of the original loss plus an importance penalty. Minimizing this objective while resampling the mask at each step would force the adversarial example to distribute the importance over all parameters, making it more robust to different parameters and thus more transferable.
Practical Implementation with DropConnect
@@ 3.2 Alleviating Over-reliance via Random Parameter Pruning
The above random parameter pruning method is similar to DropConnect method [44] in training neural networks. We notice that DropConnect is mainly effective in terms of linear layers . We thus apply DropConnect to the weight and bias parameters of linear layers as well as the transformation parameters of normalization layers. Both types of layers are widely used in mainstream architectures including Transformer [43]. Empirically, our ablation study in Section 4 validates the effectiveness of this choice, compared with convolutional layers.
@@
We now present our attack method, Random Parameter Pruning Attack (RaPA), as summarized in Algorithm 1. Take linear layer for example. We perform independent random masking onto the weight and bias (if present) parameters using Bernoulli sampling. Specifically, for surrogate model $f$, let $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ denote the weight matrix and $b\in\mathbb{R}^{d_{\text{out}}}$ the bias vector associated with a linear layer. Here $d_{\text{in}}$ and $d_{\text{out}}$ represent the input and output dimensions, respectively. The corresponding masks for the weight matrix and bias are
$\mathcal{M}_{w}\sim\text{Bernoulli}(1-p_{w}),\mathcal{M}_{b}\sim\text{Bernoulli}(1-p_{b}),$ (5)
@@
$W_{\mathcal{M}}=\mathcal{M}_{w}\odot W,~~b_{\mathcal{M}}=\mathcal{M}_{b}\odot b,$ (6)
@@
The random masks $\mathcal{M}_{w}$ and $\mathcal{M}_{b}$ in Eq. (5) are re-generated for each inference, producing diverse variants of the surrogate model with different parameters if we conduct multiple inferences at an iteration. In addition, RaPA can be naturally integrated with existing input transformation and gradient stabilization methods for crafting adversarial examples, as shown in Lines 5 and 8 in Algorithm 1.
Algorithm 1 Random Parameter Pruning Attack(RaPA)
@@
6:   Obtain modified model $f_{\mathcal{M}}$ by applying RaPA to each layer in $\mathbb{L}$ according to Eqs. (5) and (6).
@@
Method DI RDI SI Admix ODI BSR CFM RaPA
All Layer Average 0.32 0.30 0.21 0.12 0.33 0.25 0.19 0.08
Conv Layer 0.11 0.07 0.03 0.01 0.12 0.05 0.03 0.00
Norm Layer 0.51 0.52 0.37 0.22 0.53 0.44 0.34 0.15
Linear Layer 1.00 0.86 0.59 0.18 1.00 0.75 0.55 0.13
Table 2: Gini coefficients of parameter importance across different layers and methods. Lower values correspond to more uniform parameter importance.
Analyzing Parameter Importance with Gini Coefficient
To further verify that our random parameter pruning strategy indeed mitigates the over-reliance on a few dominant parameters, we employ the Gini coefficient to measure the distribution of parameter importance across layers. A lower Gini value indicates a more uniform distribution of importance, implying that the adversarial perturbation depends less on specific parameters and generalizes better to unseen models. The formal definition and detailed computation process of the Gini coefficient are provided in Sec. B.3. .
We compute the Gini coefficients based on the parameter importance values $\mathcal{I}(\theta_{i})$ defined in Eq. (3). The overall and layer-wise results are summarized in Table 2. As shown in Table 2, RaPA achieves the lowest Gini coefficients among all compared methods, suggesting that it effectively flattens the importance distribution and suppresses the over-concentration of sensitivity on a few parameters. This balanced importance allocation leads to improved robustness and transferability across different architectures.